Solve the equation and express its roots in simplest a + bi form:

2(x-3) = - 5 divided by x

Answer by stanbon(75887)

2(x-3) = - 5 divided by x
------
2x-6 = -5/x
2x^2 - 6x + 5 = 0
x = [6 +- sqrt(36 - 4*2*5)]/4
x = [6 +- sqrt(76)]/4
x = [6 + 2isqrt(19)]/4 or x = [6 + 2isqrt(19)]/4
x = (3/2) + (1/2)isqrt(19) or x = (3/2) - (1/2)isqrt(19)
===============================
Cheers,
Stan H.